# Position: On-Premises LLM Deployment Demands a Middle Path: Preserving Privacy Without Sacrificing Model Confidentiality

### v2

@@ 7 Conclusion
In this position paper, we advocate for a balanced approach to on-premises LLM deployment, addressing the dilemma between data privacy and model confidentiality. Building on our theorem, we propose SOLID, a simple yet effective semi-open solution that efficiently secures a few bottom layers of LLMs. We believe that our work shifts the view from skepticism to opportunity, paving a viable middle path for secure AI deployment that protects user privacy, safeguards vendor intellectual property, and preserves model customization flexibility. While our position may invite both support and critique, we hope it provides meaningful insights and fosters ongoing discourse. If this paper contributes to constructive engagement within the community, it will have fulfilled its purpose.
Impact Statement
The primary objective of this position paper is to advocate for a balanced approach to on-premises LLM deployment, ensuring both data privacy and model security. We introduce an efficient theoretically inspired semi-open deployment framework that mitigates functional extraction risks while enabling customization, providing a preliminary yet effective solution for privacy-sensitive industries. While our work primarily contributes to academic discourse, it also has broader implications for secure AI deployment in sectors where data confidentiality is critical. Ethically, we highlight the potential to simultaneously prevent data leakage, safeguard model security, and mitigate misuse risks. While we foresee no immediate societal consequences requiring specific emphasis, we acknowledge the importance of ongoing ethical considerations and the potential for future societal impacts as this interdisciplinary field evolves.
@@ Appendix A Proof of Theorem 1
In this section, we prove Theorem 1. We first revisit the our model, present several important lemmas and finally present the proof.
@@ A.2 Bounds on Different Orthogonal Components
$\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]\right|\leq(1+\beta_{D})\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|,$ (4)
where $\mathbb{I}_{n}$ is a column vector with dimensions $n\times 1$ and each element is 1, $\mathbf{X}[p]$ is the $p$-th column of the input $\mathbf{X}$, $\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]$ is the $p$-th column of the $l$-th self-attention output, the coefficient $\beta_{D}$ satisfies $10<\beta_{D}<$ 1 and it is related to the upper bound of the L2-norm of matrices $K_{l},Q_{l}$.
@@
Let $\bm{u}=\left\{\bm{u}_{l,1}=\frac{\mathbb{I}_{n}}{\sqrt{n}},\bm{u}_{l,2},\ldots,\bm{u}_{l,n}\right\}$ denote the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Assume $\sigma_{l,1},\sigma_{l,2},\ldots,\sigma_{l,n}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$ and $-1<\sigma_{l,n}<\beta_{D}$ for any $l,n$. Thus we have
$\bm{v}^{\top}\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]$ $=\bm{v}^{\top}\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\mathbf{X}[p]$ (5a)
$=\bm{v}^{\top}\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\sum_{k=1}^{n}\alpha_{pk}\bm{u}_{l,k}$ (5b)
$=\bm{v}^{\top}\sum_{k=1}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}$ (5c)
$\leq\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{v}^{\top}{u}_{l,k}\right|$ (5d)
$=\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}\right\|_{2}$ (5e)
$=\left[\sum_{k=2}^{n}\alpha^{2}_{pk}(1+\sigma_{l,k})^{2}\right]^{1/2}$ (5f)
$\leq(1+\beta_{D})\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|,$ (5g)
@@
$\beta_{D}=\max_{\left\|K_{l}\right\|_{2}\leq D,\left\|Q_{l}\right\|_{2}\leq D}\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}{\left\|\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\bm{v}\right\|_{2}}<1.$
The equation (5c) is due to $\bm{u}_{l,k}$ are the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. The inequality (5e) is because when $\bm{v}=\frac{\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}}{\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k}){u}_{l,k}\right\|_{2}}$, we have the maximum value.
@@
Let $\sigma_{i1},\sigma_{i2},\ldots,\sigma_{in}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Applying the Perron–Frobenius theorem for Markov matrices (Lemmens & Nussbaum, 2012), we deduce that for the matrix $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$, there exists only one eigenvalue equal to 1, while all other eigenvalues in absolute value are strictly less than 1. Without loss of generality, we assume $\sigma_{i1}=1$, implying $\left|\sigma_{ij}\right|<1$ for $j\neq 1$. Recalling the definition of $\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)$ and considering the linear operation, we can rewrite it as follows:
@@
where (10a) is induced by substituting the equation (9) into $\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$, (10b) is due to $\mathbb{I}_{n}^{\top}\bm{u_{j}}=0$ for all $j\neq 1$, (10c) follows the fact that $\sigma_{i1}=1$
@@ A.3 Proof of Theorem 1
We first prove the following result. For simplicity of notations, we use $f(\mathbf{X})\left[p\right]$ to denote the $p$-th ($1\leq p\leq d$) column of the the distilled model $f(\mathbf{X})$, where the parameters in the $\alpha L$-th layer is replaced with the matrices $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$. We use the function $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$ to denote the mapping of the $(\alpha L)$-th layer. Then we are going to show that there exists $\alpha^{\star}=\log_{2}{\frac{2}{1+\beta_{D}}}$ and $10<\beta_{D}<$ 1 makes the following equations hold.
@@
$\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}f\left(\mathbf{X}\right)[p]\right|\leq(1+\beta)^{L}\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|.$ (13)
@@
$\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|.$ (14)
@@
$\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|\neq 0,$ (15)
@@
$\left\|\frac{f(\mathbf{X})\left[p\right]}{\left\|f(\mathbf{X})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}$ $=\left[2-\frac{2\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]}{\sqrt{n}\sqrt{\frac{(\mathbb{I}_{n}^{\top}f(\mathbf{X})[p])^{2}}{n}+(\bm{v}^{\top}f(\mathbf{X})[p])^{2})}}\right]^{1/2}$ (16a)
$=\sqrt{2}\left[1-\frac{1}{\sqrt{1+\frac{n(\bm{v}^{\top}f(\mathbf{X})[p])^{2}}{(\mathbb{I}_{n}^{\top}f(\mathbf{X})[p])^{2}}}}\right]^{1/2}$ (16b)
$\leq\sqrt{2}\left[1-\frac{1}{\sqrt{1+\frac{n(1+\beta)^{2L}|\bm{v}^{\top}\mathbf{X}[p]|^{2}}{2^{2[(1-\alpha)L-1]}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|^{2}}}}\right]^{1/2}$ (16c)
$\leq 2\sqrt{2n}\left(\frac{1+\beta}{2^{1-\alpha}}\right)^{L}\frac{\left|\bm{v}^{\top}\mathbf{X}[p]\right|}{\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|},$ (16d)
where the inequality (16c) is based on the inequality (13) and (14). The inequality (16d) is based on Lemma (3). Therefore, if $\alpha<\log_{2}{\frac{2}{1+\beta_{D}}}$ and $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|\neq 0$, then we have $\lim_{L\to\infty}\left(\frac{1+\beta_{D}}{2^{1-\alpha}}\right)^{L}=0$. Now we can consider when $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0.$ In fact, it is easy to show that this can only happens when $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ belong to certain sets making $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0,$ which corresponds to zero measure set $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ depending on the input $\mathbf{X}$. Since the input space is countable, therefore, the union $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$ are also zero-measure sets.
@@
$f(\mathbf{X}^{\star})\left[p\right]=\frac{\mathbb{I}_{n}^{\top}}{\sqrt{n}}f(\mathbf{X}^{\star})\frac{\mathbb{I}_{n}}{\sqrt{n}}+\bm{v}^{\star\top}f(\mathbf{X}^{\star})\bm{v}^{\star}+\bm{u}^{\star\top}f(\mathbf{X}^{\star})\bm{u}^{\star}.$ (18)
@@
$\left|\bm{v}^{*\top}f\left(\mathbf{X}^{\star}\right)[p]\right|=(1+\beta_{D})^{L}\left|\bm{v}^{*\top}\mathbf{X}^{\star}[p]\right|\neq 0,$ (21)

### v3

@@ 6 Conclusion
In this paper, we explore minimal secured sets to protect LLMs from query-based distillation attacks while preserving customization flexibility in on-premises deployments. We find that (1) distillation attacks targeting the secured top layer can successfully replicate the victim model, and (2) both the placement and number of secured layers introduce a security-customization trade-off. Based on these insights, we propose SOLID, a theoretically inspired deployment that optimizes this trade-off. Through extensive experiments, we show that SOLID balances security and customization effectively, outperforming baseline deployments, though it also has certain limitations.
Limitations
While our method effectively defends against distillation attacks and preserves model customization, it does not address other adversarial attacks in the black-box setting, such as membership inference attacks (MIA), as demonstrated in the Appendix B.10. To the best of our knowledge, this is the first work to explore a semi-open deployment framework for LLMs. However, the current algorithm still performs identification at the layer level and does not delve into the impact of different submodules within decoder layers on model security. Furthermore, our proposed metric exhibits reduced effectiveness when applied to smaller models, as discussed in Section 4.5. In future work, we aim to address these limitations by enhancing both the algorithm and the evaluation metric to improve the overall effectiveness of SOLID.
Acknowledgements
We are deeply grateful to the anonymous reviewers for their insightful comments and constructive suggestions, which significantly improved this work. We also acknowledge the support from the National Natural Science Foundation of China (No. 62306179, No.62472247 and No. 12326608); the National Key Research and Development Program of China(2024YFB3310000); Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone Project (No.HZQSWS-KCCYB-2024016); Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001) and the Scientific Research Program of Shanghai Municipal Science and Technology Commission (24BC3200100).
@@ Appendix A Proof of Theorem 1
In this section, we prove Theorem 1. We first revisit the our model, present several important lemmas and finally present the proof. Additional explanatory remarks are included in Appendix A.5.
@@ A.2 Bounds on Different Orthogonal Components
$\begin{split}\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}&\left|\bm{v}^{\top}\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]\right|\\ &\leq(1+\beta_{D})\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|\end{split}$ (4)
where $\mathbb{I}_{n}$ is a column vector with dimensions $n\times 1$ and each element is 1, $\mathbf{X}[p]$ is the $p$-th column of the input $\mathbf{X}$, $\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]$ is the $p$-th column of the $l$-th self-attention output, the coefficient $\beta_{D}$ satisfies $0<\beta_{D}<1$ and it is related to the upper bound of the L2-norm of matrices $K_{l},Q_{l}$.
@@
Let $\bm{u}=\left\{\bm{u}_{l,1}=\frac{\mathbb{I}_{n}}{\sqrt{n}},\bm{u}_{l,2},\ldots,\bm{u}_{l,n}\right\}$ denote the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Assume $\sigma_{l,1},\sigma_{l,2},\ldots,\sigma_{l,n}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$ and $1<\sigma_{l,n}<\beta_{D}$ for any $l,n$. Thus we have
$\bm{v}^{\top}\varphi_{l}\left(\mathbf{X};K_{l},Q_{l}\right)[p]$ (5a)
$=\bm{v}^{\top}\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\mathbf{X}[p]$ (5b)
$=\bm{v}^{\top}\left[\mathbf{I}_{n}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\right]\sum_{k=1}^{n}\alpha_{pk}\bm{u}_{l,k}$ (5c)
$=\bm{v}^{\top}\sum_{k=1}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}$ (5d)
$\leq\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{v}^{\top}{u}_{l,k}\right|$ (5e)
$=\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}\right\|_{2}$ (5f)
$=\left[\sum_{k=2}^{n}\alpha^{2}_{pk}(1+\sigma_{l,k})^{2}\right]^{1/2}$ (5g)
$\leq(1+\beta_{D})\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|,$ (5h)
@@
$\beta_{D}$ $=\max_{\begin{subarray}{c}\left\|K_{l}\right\|_{2}\leq D,\\ \left\|Q_{l}\right\|_{2}\leq D\end{subarray}}\max_{\begin{subarray}{c}\bm{v}:\left\|\bm{v}\right\|_{2}=1,\\ \bm{v}\perp\mathbb{I}_{n}\end{subarray}}\left\|\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\bm{v}\right\|_{2}$
$<1$
The equation (5d) is due to $\bm{u}_{l,k}$ are the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. The inequality (5f) is because when $\bm{v}=\frac{\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}}{\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k}){u}_{l,k}\right\|_{2}}$, we have the maximum value.
@@
Let $\sigma_{i1},\sigma_{i2},\ldots,\sigma_{in}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Applying the Perron–Frobenius theorem for Markov matrices Lemmens and Nussbaum (2012), we deduce that for the matrix $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$, there exists only one eigenvalue equal to 1, while all other eigenvalues in absolute value are strictly less than 1. Without loss of generality, we assume $\sigma_{i1}=1$, implying $\left|\sigma_{ij}\right|<1$ for $j\neq 1$. Recalling the definition of $\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)$ and considering the linear operation, we can rewrite it as follows:
@@
where (10a) is induced by substituting the equation (9) into $\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$, (10b) is due to $\mathbb{I}_{n}^{\top}\bm{u_{j}}=0$ for all $j\neq 1$, (10c) follows the fact that $\sigma_{i1}=1$ .
@@ A.3 Proof of Theorem 1
We first prove the following result. For simplicity of notations, we use $f(\mathbf{X})\left[p\right]$ to denote the $p$-th ($1\leq p\leq d$) column of the the distilled model $f(\mathbf{X})$, where the parameters in the $\alpha L$-th layer is replaced with the matrices $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$. We use the function $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$ to denote the mapping of the $(\alpha L)$-th layer. Then we are going to show that there exists $\alpha^{\star}=\log_{2}{\frac{2}{1+\beta_{D}}}$ and $0<\beta_{D}<1$ makes the following equations hold.
@@
$\begin{split}&\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\;\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}f\left(\mathbf{X}\right)[p]\right|\\ &\leq\;(1+\beta)^{L}\max_{\bm{v}:\left\|\bm{v}\right\|_{2}=1,\;\bm{v}\perp\mathbb{I}_{n}}\left|\bm{v}^{\top}\mathbf{X}[p]\right|.\end{split}$ (13)
@@
$\begin{split}&\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|\\ &=\,2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|.\end{split}$ (14)
@@
$\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|\neq 0,$ (15)
@@
$\left\|\frac{f(\mathbf{X})\left[p\right]}{\left\|f(\mathbf{X})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}$ (16a)
$=\left[2-\frac{2\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]}{\sqrt{n}\sqrt{\frac{(\mathbb{I}_{n}^{\top}f(\mathbf{X})[p])^{2}}{n}+(\bm{v}^{\top}f(\mathbf{X})[p])^{2})}}\right]^{1/2}$ (16b)
$=\sqrt{2}\left[1-\frac{1}{\sqrt{1+\frac{n(\bm{v}^{\top}f(\mathbf{X})[p])^{2}}{(\mathbb{I}_{n}^{\top}f(\mathbf{X})[p])^{2}}}}\right]^{1/2}$ (16c)
$\leq\sqrt{2}\left[1-\frac{1}{\sqrt{1+\frac{n(1+\beta)^{2L}|\bm{v}^{\top}\mathbf{X}[p]|^{2}}{2^{2[(1-\alpha)L-1]}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|^{2}}}}\right]^{1/2}$ (16d)
$\leq 2\sqrt{2n}\left(\frac{1+\beta}{2^{1-\alpha}}\right)^{L}\frac{\left|\bm{v}^{\top}\mathbf{X}[p]\right|}{\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X})[p]\right|},$ (16e)
where the inequality (16d) is based on the inequality (13) and (14). The inequality (16e) is based on Lemma (3). Therefore, if $\alpha<\log_{2}{\frac{2}{1+\beta_{D}}}$ and $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|\neq 0$, then we have $\lim_{L\to\infty}\left(\frac{1+\beta_{D}}{2^{1-\alpha}}\right)^{L}=0$. Now we can consider when $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0.$ In fact, it is easy to show that this can only happens when $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ belong to certain sets making $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0,$ which corresponds to zero measure set $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ depending on the input $\mathbf{X}$. Since the input space is countable, therefore, the union $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$ are also zero-measure sets.
@@
$\begin{split}f(\mathbf{X}^{\star})\left[p\right]&=\frac{\mathbb{I}_{n}^{\top}}{\sqrt{n}}f(\mathbf{X}^{\star})\frac{\mathbb{I}_{n}}{\sqrt{n}}\\ &+\bm{v}^{\star\top}f(\mathbf{X}^{\star})\bm{v}^{\star}+\bm{u}^{\star\top}f(\mathbf{X}^{\star})\bm{u}^{\star}.\end{split}$ (18)
@@
$\left|\bm{v}^{*\top}f\left(\mathbf{X}^{\star}\right)[p]\right|=(1+\beta_{D})^{L}\left|\bm{v}^{*\top}\mathbf{X}^{\star}[p]\right|\neq 0.$ (21)